# STAPO: Stabilizing Reinforcement Learning for LLMs by Silencing Rare Spurious Tokens

### v1

@@ 5.5.2 Word Cloud Visualization
Figure 6: Analysis of Spurious Tokens: (a) indicates the ratio of spurious tokens during training; (b) and (c) present the word clouds of frequently masked spurious tokens and retained normal tokens
@@ 6 Conclusion
In this work, we investigate the instability of RL from a token-level perspective and identify a pathological update regime driven by the interaction between token probability, policy entropy, and advantage sign. We show that a negligible fraction of spurious tokens, defined as rare and uninformative tokens that coincidentally co-occur with correct outcomes, can receive disproportionately amplified gradients and exert an outsized influence on optimization dynamics. Based on this analysis, we propose STAPO, a simple yet effective method that suppresses approxim ately $0.01\%$ of destructive token updates and renormalizes the loss over valid tokens. Empirical results across six mathematical reasoning benchmarks demonstrate that this minimal intervention substantially stabilizes policy entropy and consistently improves reasoning performance.

### v2

@@ 5.5.2 Word Cloud Visualization
Figure 6: Analysis of Spurious Tokens. (a) indicates the ratio of spurious tokens during training; (b) and (c) present the word clouds of frequently masked spurious tokens and retained normal tokens
@@ 6 Conclusion
In this work, we investigate the instability of RL from a token-level perspective and identify a pathological update regime driven by the interaction between token probability, policy entropy, and advantage sign. We show that a negligible fraction of spurious tokens, defined as rare and uninformative tokens that coincidentally co-occur with correct outcomes, can receive disproportionately amplified gradients and exert an outsized influence on optimization dynamics. To address this instability, we equip the training process with S2T mechanism, that efficiently detects spurious tokens using characteristic signals including low probability, low entropy, and positive advantage, and curtails their gradient influence during optimization. Embedded within a group-based objective, we propose Spurious-Token-Aware Policy Optimization (STAPO), a framework designed to enhance the stability and effectiveness of large-scale model refinement. Empirical results across six mathematical reasoning benchmarks demonstrate that this minimal intervention substantially stabilizes policy entropy and consistently improves reasoning performance.